in one city 52.6% of adults are female . 9.5% of adults are left handed and 5.3 % are left handed female p( f or l ) . round approximation to the three decimal places


1
Expert's answer
2021-05-10T13:45:38-0400

The events for the question are as follows:

"F=Event\:where\:the\:adult\:is\:female"

"L=Event\:where\:the\:adult\:is\:left\:handed"

"FL\:=\:Event\:where\:the\:adult\:is\:\:and\:left\:handed\:female\:"


To find the probability of at least one of the two joint events we add the probabilities of these two events excluding the probability of their co-occurrence.

"P\left(ForL\right)=P\left(F\right)+P\left(L\right)-P\left(FL\right)"

Where:

"P\left(F\right)" "=Probability\:of\:event\:F\:=\:0.526"

"P\left(L\right)""=\:Probability\:of\:event\:L\:=\:0.095"

"P\left(FL\right)" "=Probability\:of\:event\:F\:and\:L\:=\:0.053"

"P\left(ForL\right)=0.526+0.095-0.053\:"

"P\left(ForL\right)=0.568\:"